OpenStudy (anonymous):

6x+x(x-1)=0, what are the solutions? x=?

OpenStudy (anonymous):

I got x=0, -5

OpenStudy (anonymous):

6x+x^2-x=0 x^2 +5x =0 x(x+5)=0 x=0 x=-5

OpenStudy (anonymous):

thank